# Boosting Omnidirectional Stereo Matching with a Pre-trained Depth Foundation Model

### v2

@@ I INTRODUCTION
In this paper, we introduce DFI-OmniStereo (Depth Foundation Model-based Iterative Omnidirectional Stereo Matching), a novel omnidirectional stereo matching method that integrates a pre-trained monocular depth foundation model into an iterative optimization-based stereo matching architecture, as illustrated in Figure 1. Our approach follows a two-stage training strategy in which (i) the stereo matching head learns to adapt to the new feature space and camera setup while keeping the foundation model fixed, and then (ii) the foundation model’s decoder is unfrozen to be fine-tuned using a scale-invariant loss without foregoing its generalization capabilities. We conduct extensive experiments to compare DFI-OmniStereo with other omnidirectional stereo matching methods, perform detailed analyses, and evaluate its sample efficiency as well as its generalization capabilities.
Specifically, our contributions are as follows: (1) We leverage a large-scale pre-trained monocular depth foundation model as a feature extractor integrated into an iterative optimization-based stereo matching architecture. (2) We design a two-stage training strategy to adapt monocular foundation model features to the omnidirectional stereo matching setup. We partially fine-tune the foundation model, and employ a scale-invariant error in log space (SILog loss) for stereo matching. (3) Our method demonstrates state-of-the-art results on the Helvipad [8] dataset, a challenging real-world benchmark for omnidirectional stereo matching. In addition, DFI-OmniStereo shows generalization capabilities to other datasets and high sample efficiency.
@@ II RELATED WORK
A disparity map can be estimated from two images and then deterministically converted to depth assuming a calibrated stereo setup. Early deep learning methods [12] use Convolutional Neural Networks (CNNs) for matching cost calculation or end-to-end disparity estimation [13]. More recent 3D networks [14, 15, 16, 17] introduce 4D cost volumes via feature concatenation and employ a 3D encoder-decoder design for context aggregation. Many recent works in deep stereo matching follow two paradigms: Vision Transformer [18] (ViT)-based methods and iterative optimization-based methods [19]. GMStereo [20] and CroCo-Stereo [21] propose a unified transformer-based architecture for optical flow and stereo. RAFT-Stereo [22] presents an iterative multi-resolution approach that refines disparity maps using a multi-level 3D correlation volume and a learned context encoding via convolutional GRUs [23]. Following RAFT-Stereo, many subsequent architectures [24, 25, 26, 27] propose an iterative refinement of the disparity. In particular, IGEV-Stereo [28] constructs a 3D cost volume that fuses local cues from a multi-level 3D correlation volume with global context from a regularized 4D cost volume to iteratively refine disparity estimates. Vankadar et al. [29] and ViTAStereo [30] use features from the vision foundation models DINO [31] and DINOv2 [32] for stereo matching. Our approach extends this family by incorporating a large-scale pre-trained depth foundation model as feature extractor to consider the strong correlation between relative depth and disparity. Unlike concurrent work,111 Note that, independently and concurrently with our work, Wen et al. [33] and Cheng et al. [34] adopt a conceptually similar methodology to DFI-OmniStereo, applying it to conventional stereo matching. We also fine-tune the decoder of a foundation model to (omnidirectional) stereo matching.
@@ IV EXPERIMENTS
In this section, we compare DFI-OmniStereo against existing stereo-matching methods on real-world data and provide insights into its accuracy. We further study the model’s sample-efficiency, in particular in low training data regimes, and its generalization capabilities.
@@ IV-D Analyzing DFI-OmniStereo
IV-E Sample-efficient Learning
Collecting labeled real-world data is expensive. Having methods that can learn from a small amount of samples is essential for real-world applications. Figure 4 shows that DFI-OmniStereo already achieves a lower disparity MAE than 360-IGEV-Stereo (with $100\%$ training data) when only $5\%$ of the training data, randomly sampled, are available to our model. This highlights how leveraging large-scale pre-training from a foundation model significantly reduces the need for task-specific data.
@@ IV-E Sample-efficient Learning
Figure 4: Sample-efficient learning analysis using DFI-OmniStereo on the Helvipad dataset [8]. The training data for our method is a randomly sampled subset. 360-IGEV-Stereo [8] is visualized as the dashed line using $100\%$ of the training data for comparison.

### v3

@@ I INTRODUCTION
In this paper, we introduce DFI-OmniStereo (Depth Foundation Model-based Iterative Omnidirectional Stereo Matching), a novel omnidirectional stereo matching method that integrates a pre-trained monocular depth foundation model into an iterative optimization-based stereo matching architecture, as illustrated in Figure 1. Our approach follows a two-stage training strategy in which (i) the stereo matching head learns to adapt to the new feature space and camera setup while keeping the foundation model fixed, and then (ii) the foundation model’s decoder is unfrozen to be fine-tuned using a scale-invariant loss without foregoing its generalization capabilities. We conduct extensive experiments to compare DFI-OmniStereo with other omnidirectional stereo matching methods, perform detailed analyses, and evaluate its training sample efficiency as well as its generalization capabilities.
Specifically, our contributions are as follows: (1) We leverage a large-scale pre-trained monocular depth foundation model as a feature extractor integrated into an iterative optimization-based stereo matching architecture. (2) We design a two-stage training strategy to adapt monocular foundation model features to the omnidirectional stereo matching setup. We partially fine-tune the foundation model, and employ a scale-invariant error in log space (SILog loss) for stereo matching. (3) Our method demonstrates state-of-the-art results on the Helvipad [8] dataset, a challenging real-world benchmark for omnidirectional stereo matching. In addition, DFI-OmniStereo shows generalization capabilities to other datasets and high training sample efficiency.
@@ II RELATED WORK
A disparity map can be estimated from two images and then deterministically converted to depth assuming a calibrated stereo setup. Early deep learning methods [12] use Convolutional Neural Networks (CNNs) for matching cost calculation or end-to-end disparity estimation [13]. More recent 3D networks [14, 15, 16, 17] introduce 4D cost volumes via feature concatenation and employ a 3D encoder-decoder design for context aggregation. Many recent works in deep stereo matching follow two paradigms: Vision Transformer [18] (ViT)-based methods and iterative optimization-based methods [19]. GMStereo [20] and CroCo-Stereo [21] propose a unified transformer-based architecture for optical flow and stereo. RAFT-Stereo [22] presents an iterative multi-resolution approach that refines disparity maps using a multi-level 3D correlation volume and a learned context encoding via convolutional GRUs [23]. Following RAFT-Stereo, many subsequent architectures [24, 25, 26, 27] propose an iterative refinement of the disparity. In particular, IGEV-Stereo [28] constructs a 3D cost volume that fuses local cues from a multi-level 3D correlation volume with global context from a regularized 4D cost volume to iteratively refine disparity estimates. Vankadar et al. [29] and ViTAStereo [30] use features from the vision foundation models DINO [31] and DINOv2 [32] for stereo matching. Our approach extends this family by incorporating a large-scale pre-trained depth foundation model as feature extractor to consider the strong correlation between relative depth and disparity. [^1] Note that, independently and concurrently with our work, Wen et al. [33] and Cheng et al. [34] adopt a conceptually similar methodology to DFI-OmniStereo, applying it to conventional stereo matching. We also fine-tune the decoder of a foundation model to (omnidirectional) stereo matching.
@@ IV EXPERIMENTS
In this section, we compare DFI-OmniStereo against existing stereo-matching methods on real-world data and provide insights into its accuracy. We further study the model’s training sample-efficiency, in particular in low training data regimes, and its generalization capabilities.
@@ IV-D Analyzing DFI-OmniStereo
IV-E Training Sample-efficient Learning
Collecting labeled real-world data is expensive. Having methods that can learn from a small amount of training samples is essential for real-world applications. Figure 4 shows that DFI-OmniStereo already achieves a lower disparity MAE than 360-IGEV-Stereo (with $100\%$ training data) when only $5\%$ of the training data, randomly sampled, are available to our model. This highlights how leveraging large-scale pre-training from a foundation model significantly reduces the need for task-specific data.
@@ IV-E Training Sample-efficient Learning
Figure 4: Training sample-efficient learning analysis using DFI-OmniStereo on the Helvipad dataset [8]. The training data for our method is a randomly sampled subset. 360-IGEV-Stereo [8] is visualized as the dashed line using $100\%$ of the training data for comparison.